Find the exact solution of y=x+2
y=x^2

Answer by checkley77(12844)

y=x+2
y=x^2 subtract.
--------
0=-x^2+x+2
x^2-x-2=0
(x-2)(x+1)=0
x-2=0
x=2 ans.
x+1=0
x=-1 ans.